Commit 48169ad8 authored by Mei's avatar Mei
Browse files

Add new file

parent 5d2c0369
# Work discussion 10-10-2017
Present: Pieter Neerincx (PN), Matthijs Moed (MM), Leon Mei (LM), Peter van 't Hof (PH), Mark Santcroos (MS)
## COmanage
- Confusion was identified at the DTL programmer meeting. Due to maintenance at HPC cloud, Pieter/Mark are testing more testing this week. Dynamic IP for VM is not available anymore. Pieter/Mark will add HPC cloud IP range to their local firewall to allow the connection.
- Mark: Data seems lost at the VM because the VM was not properly shutdown. Gerben is working on getting a new VM started by tomorrow.
- For production CoManage VM, we need to look into a different support/organization.
- Gerben will follow up with Leon/Pieter/Mark based on work plan layed down by Matthijs. More IdPs are needed to support a demo at Health-RI meeting.
## GoNL re-analysis
- New genotype calling is done now. PH is working on QC and making draft deliverable of VCF call set.
- PN removed the wrong lanes from sample 58C from all GoNL sample sheets.
## Demo on Dec 8th.
WP5 MT agreed 3 demos for the pipeline part.
- GoNL new result (describing effort, speed, bammetrics, VCFstat plot, etc). Peter is working on the QC.
- Ansible script to deploy BIOS (biopet) and GoNL (Molgenis) pipelines. BIOS is ready and GoNL pipeline is pending. We will rename the repos. PN will make a first draft release of DNA pipeline script tomorrow.
- PIM viewer for biopet pipeline. Peter is working with Thomas from WP3 on this.
## Ansible
PN is polishing his ansible scripts and PH shared some suggestions. More questions can be sent to PH.
\ No newline at end of file
Markdown is supported
0% or .
You are about to add 0 people to the discussion. Proceed with caution.
Finish editing this message first!
Please register or to comment